# pneudots.bib

### v1

@@ Abstract
Soft pneumatic actuators have a wide range of applications, including providing haptic feedback embedded in smart garments. Here we investigate actuators fabricated from thermoplastic coated textiles. We measure the effects of fabrication parameters on the robustness and airtightness of small, round pneumatic pouch actuators made from heat-sealed thermoplastic polyurethane-coated nylon, which we call PneuDots. We determine the optimal temperature, time, and pressure for heat-pressing of the textile to create strong bonds and identify the most effective glue to create an airtight seal at the inlet. Compared to elastomeric pneumatic actuators, PneuDots reduce the thickness of the actuator by 96.4% and the mass by 57.2%, increasing their wearability while maintaining a strong force output. We evaluated the force output of the actuators, along with their performance over time. In a blocked force test, PneuDot maximum force transmission was 36.1 N, which is 95.3% of the peak force output of an elastomeric pneumatic actuator with the same diameter and pressure. Cyclical testing showed that PneuDots had more stable behavior over time. These results provide best practices for fabrication and indicate the feasibility of textile pneumatic actuators for future wearable applications.
@@ I Introduction
Figure 1: A. Thermoplastic pouch pneumatic actuators in a vest [Nunez2022AActuation]. B. 3D-printed elastomeric pneumatic actuators in a sleeve [duPasquier2023ASleeve]. C. PneuDots textile pneumatic actuators, which are developed and tested in this paper. We show an enlarged view of the 3D-printed inlet.
@@
Pneumatic actuators are preferred over other types of soft robotic actuation due to their high strength-to-weight ratio [Jang2023DesignActuation][Niiyama2014PouchRobotics]. Textile and film-based actuators are slim and lightweight pouches which can conform to the body and produce large force outputs [ONeill2017ATesting]. A thin and flexible pouch inflates into a large and stiff volume; this transformation is particularly attractive for wearable devices, and specifically haptic devices.
@@
Pneumatic haptic devices have emerged as a promising alternative to vibration-based systems, offering a broader and more nuanced range of haptic feedback [duPasquierHaptiknit:Haptics]. Using pneumatic pouches in contact with the skin, these devices mimic the feeling of human touch [Talhan2024SoftTouch], provide situational awareness [Raitor2017WRAP:Guidance][Rognon2019SoftDrone], and enhance immersion in virtual reality [Delazio2018ForceExperiences]. Pneumatic actuators have been incorporated into bracelets [Pohl2017Squeezeback:Notifications][Jumet2022ADevices], sleeves [Zhu2020PneuSleeve:Sleeve][Liu2021ThermoCaress:Stimulation][Choi2023DevelopmentActuator][Jumet2023FluidicallyTextiles], fingertip bands [VanBeek2024ValidationFeedback][Frediani2020TactileFingertip], leg bands [Endow2021Compressables:Interfaces][Fan2009PilotAmputee], and vests [Nunez2022AActuation][Delazio2018ForceExperiences][Jadhav2023ScalableActuators][Kang2023PneumaticSimulation] to provide a wide variety of cues and sensations, ranging from comforting human touch to collisions in virtual reality games.
@@
Many pneumatic actuators for haptic devices are made of thermoplastic films (polyethylene (PE), polyurethane (TPU), PVC, or similar) [Niiyama2014PouchRobotics][Niiyama2015PouchDesign][Sanchez2019DevelopmentInteraction][Yamaoka2018AccordionFab:Sheets][Connolly2019Sew-freeRobots][Do2021Macro-MiniDisplays][Wu2019WearableArm], cast elastomers [Choi2023DevelopmentActuator][Endow2021Compressables:Interfaces][Kang2023PneumaticSimulation], or coated textiles [Jumet2023FluidicallyTextiles][Ou2016AeroMorphDesign]. Thermoplastic film tubes are inexpensive and straightforward to adapt into rectangular pouches [Nunez2022AActuation][Do2021Macro-MiniDisplays]. They can be embedded into wearable devices, such as the vest shown in Fig. 1A, thanks to their thin profile. Cast elastomeric pneumatic actuators are considerably more bulky, but can stretch to inflate further than inextensible materials [Choi2023DevelopmentActuator][Endow2021Compressables:Interfaces]. 3D-printed elastomeric pneumatic actuators can be fabricated with complex custom geometries suited to specific applications or fitting specific regions of the body.
@@
Many pneumatic devices require off-board air supplies in the form of heavy pumps or compressed air to achieve high bandwidth [Jumet2023FluidicallyTextiles]. This limits their application in the home environment and as mobile devices. To make lightweight portable wearable devices, actuators should be low-volume and airtight to allow for operation by small, portable pneumatic sources, such as on-board pumps. In [duPasquierHaptiknit:Haptics], we used 3D-printed elastomeric pneumatic actuators, shown in Fig. 1B, that could be operated by a customized version of the untethered FlowIO system by Shtarbanov et al. [Shtarbanov2021FlowIORobotics]. These elastomeric pneumatic actuators were capable of generating up to 37.9 N of output force when the 491 mm2 contact area was inflated to 230 kPa absolute pressure, but are bulkier than film or textile actuators.
We present modular textile pneumatic haptic actuators that are lightweight, low-profile, comfortable, and flexible, which we call PneuDots (Fig. 1C). We focus on refining the manufacturing process to ensure strong, airtight seals to produce durable actuators. We measure the performance of the actuators and compare them to the elastomeric 3D-printed actuators proposed by du Pasquier et al. [duPasquierHaptiknit:Haptics]. This paper is structured as follows. First, we present the manufacturing method for PneuDots and examine the results of varying parameters in the manufacturing process. Then, we evaluate the performance of the textile pneumatic actuators and compare it against previous work. Finally, we contextualize the importance of this work in wearable haptic devices.
@@ II-A Fabrication
Figure 2: The PneuDots manufacturing process. A. the textile sheets and the aluminum molds. B. The heat press and its settings. C. The 3D-printed inlet and bonding adhesive.
We manufacture our textile pneumatic actuators using TPU-coated 70-denier ripstop nylon (Quest Outfitters #1056). Two pieces of nylon are placed in a heat press (DabPress Commercial Rosin Press) with the TPU-coated sides facing inwards. Above and below the fabric, we place 3/16” aluminum molds, as shown in Fig. 2A, which are in contact with the two heated plates of the press. These molds were waterjet cut to the desired geometry. Hexagonal pins ensure that the molds are aligned. Because the heat conductivity of the aluminum is higher than that of air, only the regions of the fabric that are in contact with the molds become sealed. The heat press is set to a target pressure, and the sample is kept in the press for a designated amount of time. After that time, the molds and sample are removed, the sample is set out flat to cool, and the molds are cooled with water to ensure a consistent starting temperature for every fabrication cycle. We discuss below in Section II-C how we selected the ideal temperature, pressure, and time.
@@ II-B Parameter Evaluation
The fabrication parameters which we investigate are temperature and pressure of the heat press, duration the sample was pressed, and the type of glue used to attach the inlets. Preliminary testing indicated that heat press temperatures exceeding 200 °F are required to fully bond the TPU-coated nylon, but the relationship between higher temperatures and bond quality has not yet been investigated. We test temperatures between 220 °F and 280 °F at 20 °F intervals for a total of 4 temperatures. We further test samples at three pressures, 2.5, 5, and 7.5 MPa, and pressing durations from 30 to 120 s at 30 s intervals for a total of 4 pressing durations.
@@
Figure 3: Effects of varying the manufacturing A. temperature, B. time, C. pressure, and D. adhesive on the peel strength and airtightness, with the best-performing samples shown in black.
@@
Condition 5 MPa / 260 °F / 90 s 2.5 MPa 7.5 MPa 220 °F 240 °F 280 °F 30 s 60 s 120 s
Standard Deviation (N) 14.89 9.43 11.02 16.84 8.28 37.41 19.28 18.48 28.97
Peak Force (N) 176.08 161.75 148.72 68.83 171.27 189.22 155.79 154.97 160.45
Peel Strength (N/mm) 17.82 17.19 18.74 7.49 19.24 24.49 15.91 20.29 22.60
Critical Displacement (mm) 9.06 7.03 8.47 3.81 7.20 8.13 9.31 7.87 6.09
@@ II-C Results and Discussion
We choose to use the Gear Aid adhesive to fabricate the PneuDots textile pneumatic actuators because of its airtightness, its low variability, its economical packaging, and its smoothness and flexibility when dried.
@@ III-A Methods
Using the fabrication parameters found above, we fabricate PneuDots, our textile pneumatic actuators. We assess the performance of these actuators in force transmission and in cyclical testing. We choose these tests to evaluate the ability to produce normal force, to compare to similar pneumatic actuators [duPasquierHaptiknit:Haptics], and to ensure that the chosen fabrication method is robust enough to be used for wearable haptic devices, in which each actuator will undergo thousands of cycles. The force characterization tests (n=5) are conducted following a procedure similar to that outlined in [duPasquierHaptiknit:Haptics] and adapted for textile pneumatic actuators. As shown in Fig. 4A, the actuator is placed on a 3D-printed platform with a hole in it for the actuator to expand through. When inflated, the actuator presses against a force sensor (ATI Nano17 6-axis Force/Torque Sensor) below. Above the actuator, an acrylic plate prevents the actuator from expanding away from the force sensor. The platform, force sensor, and acrylic plate are secured to a 3D printed frame. The actuator is inflated by the pneumatic controller proposed by du Pasquier et al. [duPasquierHaptiknit:Haptics]. During the test, the actuator is inflated to a series of absolute pressures from 125 to 230 kPa at 15 kPa intervals (a total of 8 peaks) and vacuumed to atmospheric pressure in between each pressure. The force is recorded at 100 Hz by the force sensor.
Cyclical testing (n=4) was performed with the same frame and load cell setup. Actuators are inflated to three target pressures, 150, 175, and 200 kPa, for 400 cycles. During each cycle, the actuator is inflated to the target pressure, held for one second, vacuumed to atmospheric pressure, and held for one second. Again, the force sensor recorded the output force at a frequency of 100 Hz. For each cycle, one peak force value was extracted from the data.
@@ III-B Results
Compared to the 3D-printed elastomeric pneumatic actuators from [duPasquierHaptiknit:Haptics], PneuDots have higher consistency in the cycling tests. At 150 and 200 kPa, the output force of the elastomeric pneumatic actuators drifted by a few Newtons over time. This is likely caused by the elastomeric actuators stretching and deforming with repeated use.
@@ IV Discussion and Conclusions
We investigated the effects of several fabrication parameters on the process of creating heat-sealed textile pneumatic actuators (PneuDots): temperature, pressure, and pressing duration. We conclude that the best parameters for fabricating PneuDots are to heat press the TPU-coated nylon at 260 °F for 90 s under 5 MPa of pressure. These parameters are specific to our setup. Different results may be obtained by varying factors such as the type or thickness of the textile or by using a heat press with a single heated plate instead of two. Independent of exact values, we show that temperature, time, and pressure can have significant effects on the peel strength, peak force, and critical displacement of heat-sealed textiles, and outline a benchmarking method to evaluate the best fabrication parameters for any given setup. We also find that Gear Aid adhesive creates the best bond to our materials, and we can conclude more broadly that key features in the choice of an adhesive for textile pneumatic actuators are the following: repeatability, packaging, and post-cure adhesion, smoothness, and pliability.
@@
In testing, our actuators produce up to 36.1 N at 230 kPa absolute pressure when placed in a blocked force testing setup, which is 95.3% of the force produced by elastomeric pneumatic actuators in a similar setup [duPasquierHaptiknit:Haptics]. In contrast, textile pneumatic actuators achieve a 96.4% reduction in thickness when uninflated, decreasing the profile from 8.3 mm to 0.3 mm. The textile pneumatic actuators also achieve a 57.2% reduction in mass, from 4.73 g to 2.03 g. This allows our textile pneumatic actuators to be incorporated into lower-profile wearable devices without significantly compromising performance. Additionally, the textile actuators demonstrated more consistent performance over time during cycling tests, suggesting greater reliability for long-term use.
@@
Jumet et al. pioneered similar textile pneumatic actuators in bracelets [Jumet2022ADevices] and a sleeve [Jumet2023FluidicallyTextiles]. For these devices, a series of actuators were produced by heat-pressing two sheets of TPU-coated nylon taffeta fabric, between which a laser-cut paper layer prevents bonding in the inflation areas. Like ours, these actuators are compliant, comfortable, and low-profile. Their 25 mm square cells offer a 625 mm2 contact area, while our 25 mm round actuators offer a 491 mm2 contact area. At 100 kPa (200 kPa absolute pressure), their actuators produce up to 9.8 N of output force in a test setup mimicking the stiffness of the human wrist. In this setup, some of the force is absorbed deforming the synthetic skin, and the actuator may expand away from the force sensor, resulting in a lower measured force than a blocked-force test. In our blocked-force setup, we measured up to 30.6 N at the same pressure, but without losses. Assuming skin deformation and outward expansion reduce the force output by no more than 68%, our force measurements indicate that our actuator geometry is more effective at converting air pressure into force. These results also highlight the value of distributed-stiffness textiles, as described in [duPasquierHaptiknit:Haptics], for controlling the deformation of embedded pneumatic actuators. Layering stiff textiles over the actuator mimics a blocked-force setup by directing actuator deformation in a single direction.
Figure 5: Potential applications for PneuDots in situational awareness, navigation, and posture feedback. (Original images from pixabay.com.)
@@
Due to to their low profile and rapid and repeatable fabrication method, PneuDots can be integrated into a variety of wearable devices, such as bracelets, sleeves, vests, collars, or pants. Their form factor allows for seamless and discrete integration into garments, increasing their wearability and user appeal. The soft textile surface of the actuators allows for comfortable nonadhesive direct-to-skin contact. Their compact inflation volume allows for operation with a portable controller equipped with an onboard pump and vacuum, thereby enhancing the potential for untethered wearability. Potential applications for garments with PneuDots, as illustrated in Fig. 5, include improving immersion in virtual reality [Delazio2018ForceExperiences][Jadhav2023ScalableActuators][Kang2023PneumaticSimulation], replicating social touch in digital environments [Talhan2024SoftTouch][Keng2008HuggySystem], assisting navigation [Raitor2017WRAP:Guidance][Jumet2023FluidicallyTextiles], or providing guidance and feedback for posture or breathing [Wu2010Torso-basedDisorders][Rotella2012HAPIInterface].

### v2

@@ Abstract
Textile pneumatic actuators can provide useful wearable haptic feedback when embedded in gloves, armbands, and other smart garments. Here we investigate actuators fabricated from thermoplastic coated textiles. We measure the effects of fabrication parameters on the robustness and airtightness of small, round pneumatic pouch actuators made from heat-sealed thermoplastic polyurethane-coated nylon. We determine the optimal temperature, time, and pressure for heat-pressing of the textile to create strong bonds and identify the most effective glue to create an airtight seal at the inlet. Compared to elastomeric pneumatic actuators, these textile pneumatic actuators reduce the thickness of the actuator by 96.4% and the mass by 57.2%, increasing their wearability while maintaining a strong force output. We evaluated the force output of the actuators, along with their performance over time. In a blocked force test, the maximum force transmission of the pneumatic textile actuators was 36.1 N, which is 95.3% of the peak force output of an elastomeric pneumatic actuator with the same diameter and pressure. Cyclical testing showed that the textile actuators had more stable behavior over time. These results provide best practices for fabrication and indicate the feasibility of textile pneumatic actuators for future wearable applications.
Index Terms:
Soft haptics, textiles, pneumatic actuators, wearable haptics
@@ I Introduction
Figure 1: A. Thermoplastic pouch pneumatic actuators in a vest [Nunez2022AActuation]. B. 3D-printed elastomeric pneumatic actuators in a sleeve [duPasquier2023ASleeve]. C. The textile pneumatic actuators tested in this paper, including an enlarged view of the 3D-printed inlet.
@@
Pneumatic actuators are often preferred over other types of soft robotic actuation due to their high strength-to-weight ratio [Jang2023DesignActuation][Niiyama2014PouchRobotics]. Textile and film-based actuators are slim and lightweight pouches, which can conform to the body and produce large force outputs [ONeill2017ATesting]. A thin and flexible pouch inflates into a large and stiff volume; this transformation is particularly attractive for wearable haptic devices.
@@
Pneumatic wearable haptic devices have emerged as a promising alternative to vibration-based systems, offering a broader and more nuanced range of haptic feedback [duPasquierHaptiknit:Haptics]. Using pneumatic pouches in contact with the skin, these devices mimic the feeling of human touch [Talhan2024SoftTouch], provide situational awareness [Raitor2017WRAP:Guidance][Rognon2019SoftDrone], and enhance immersion in virtual reality [Delazio2018ForceExperiences]. Pneumatic actuators have been incorporated into bracelets [Pohl2017Squeezeback:Notifications][Jumet2022ADevices], sleeves [Zhu2020PneuSleeve:Sleeve][Liu2021ThermoCaress:Stimulation][Choi2023DevelopmentActuator][Jumet2023FluidicallyTextiles], fingertip bands [VanBeek2024ValidationFeedback][Frediani2020TactileFingertip], leg bands [Endow2021Compressables:Interfaces][Fan2009PilotAmputee], and vests [Nunez2022AActuation][Delazio2018ForceExperiences][Jadhav2023ScalableActuators][Kang2023PneumaticSimulation] to provide a wide variety of cues and sensations, ranging from comforting human touch to collisions in virtual reality games.
@@
Many pneumatic actuators for haptic devices are made of thermoplastic films (polyethylene (PE), polyurethane (TPU), PVC, or similar) [Niiyama2014PouchRobotics][Niiyama2015PouchDesign][Sanchez2019DevelopmentInteraction][Yamaoka2018AccordionFab:Sheets][Connolly2019Sew-freeRobots][Do2021Macro-MiniDisplays][Wu2019WearableArm], cast elastomers [Choi2023DevelopmentActuator][Endow2021Compressables:Interfaces][Kang2023PneumaticSimulation], or coated textiles [Jumet2023FluidicallyTextiles][Ou2016AeroMorphDesign]. Thermoplastic film tubes are particularly attractive because they are inexpensive and straightforward to adapt into rectangular pouches [Nunez2022AActuation][Do2021Macro-MiniDisplays]. They can be embedded into wearable devices, such as the vest shown in Fig. 1A, thanks to their thin profile. Cast elastomeric pneumatic actuators are considerably more bulky, but can stretch to inflate farther than inextensible materials [Choi2023DevelopmentActuator][Endow2021Compressables:Interfaces]. 3D-printed elastomeric pneumatic actuators can be fabricated with complex custom geometries suited to specific applications or fitting specific regions of the body.
@@
Many pneumatic devices require off-board air supplies in the form of heavy pumps or compressed air to achieve high bandwidth [Jumet2023FluidicallyTextiles]. This limits their application in the home environment and as mobile devices. To make lightweight portable wearable devices, actuators should be low-volume and airtight to allow for operation by small, portable pneumatic sources, such as on-board pumps and CO2 cartridges.
The effects of size and geometry on the performance of heat-sealed thermoplastic pouch actuators have been previously explored [Jumet2023FluidicallyTextiles][Niiyama2015PouchDesign][Ou2016AeroMorphDesign], but the effects of the parameters of the fabrication processes used to produce them have not yet been characterized. Methods of sealing including heat presses and heated rollers, both manual and computerized, have been proposed and contrasted, but variables such as temperature, pressure, and contact time have not yet been investigated. We propose that these settings affect the bond strength between the layers of the actuators, impacting their reliability and performance.
We present a parameterized fabrication process for modular textile pneumatic haptic actuators that are lightweight, low-profile, comfortable, and flexible, as shown in Fig. 1C. We focus on refining the fabrication process to ensure strong, airtight seals to produce durable actuators. We measure the performance of the actuators and compare them to the elastomeric 3D-printed actuators proposed by du Pasquier et al. [duPasquierHaptiknit:Haptics] (Fig. 1B).
This paper is structured as follows: First, we present the fabrication method for the actuators and examine the results of varying parameters in the fabrication process. Then, we evaluate the performance of the textile pneumatic actuators and compare it against previous work. Finally, we contextualize the importance of this work in wearable haptic devices.
@@ II-A Fabrication
Figure 2: The textile pneumatic actuator fabrication process. A. the textile sheets and the aluminum molds. B. The heat press and its settings. C. The 3D-printed inlet and bonding adhesive.
We fabricate our textile pneumatic actuators using TPU-coated 70-denier ripstop nylon (Quest Outfitters #1056). Two pieces of nylon are placed in a heat press (DabPress Commercial Rosin Press) with the TPU-coated sides facing inwards. Above and below the fabric, we place 3/16” aluminum molds, as shown in Fig. 2A, which are in contact with the two heated plates of the press. These molds were waterjet cut to the desired geometry. Hexagonal pins ensure that the molds are aligned. Because the heat conductivity of the aluminum is higher than that of air, only the regions of the fabric that are in contact with the molds become sealed. The heat press is set to a target pressure, and the sample is kept in the press for a designated amount of time. After that time, the molds and sample are removed, the sample is set out flat to cool, and the molds are cooled with water to ensure a consistent starting temperature for every fabrication cycle. We discuss below in Section II-C how we selected the ideal temperature, pressure, and time.
@@ II-B Parameter Evaluation
The fabrication parameters that we investigate are the temperature and pressure of the heat press, the duration the sample was pressed, and the type of glue used to attach the inlets. Preliminary testing indicated that heat press temperatures exceeding 200 °F are required to fully bond the TPU-coated nylon, while temperatures above 300 °F seal areas outside of the intended contact area. To quantify the relationship between temperature and bond strength, we test temperatures between 220 °F and 280 °F at 20 °F intervals for a total of 4 temperatures. Previous literature [Jumet2022ADevices] indicates that low pressures are required to bond the textiles, so the smallest intervals on our heat press are used. We test samples at three pressures: 2.5, 5, and 7.5 MPa. In previous literature [Jumet2022ADevices][Ou2016AeroMorphDesign], pressing durations of 30 s or less were used. To determine whether longer pressing durations would produce a stronger bond, we tested from 30 to 120 s at 30 s intervals for a total of 4 pressing durations. To reduce the number of combinations to test, we designate 260 °F, 5 MPa, and 90 s as nominal values for our testing ranges and vary one parameter at a time.
@@
Figure 3: Effects of varying the fabrication A. temperature, B. time, C. pressure, and D. adhesive on the peel strength and airtightness, with the nominal values of 260 °F, 90s, and 5MPa shown in black and the critical displacement points circled.
@@
Condition Nominala 220 °F 240 °F 280 °F 30 s 60 s 120 s 2.5 MPa 7.5 MPa
Standard Deviation (N) 14.89 16.84 8.28 37.41 19.28 18.48 28.97 9.43 11.02
Peak Force (N) 176.08 68.83 171.27 189.22 155.79 154.97 160.45 161.75 148.72
Peel Strength (N/mm) 17.82 7.49 19.24 24.49 15.91 20.29 22.60 17.19 18.74
Critical Displacement (mm) 9.06 3.81 7.20 8.13 9.31 7.87 6.09 7.03 8.47
aNominal settings are 260 °F, 90s, and 5MPa.
@@ II-C Results and Discussion
We choose to use the Gear Aid adhesive to fabricate the textile pneumatic actuators because of its airtightness, its low variability, its economical packaging, and its smoothness and flexibility when dried.
@@ III-A Methods
Using the fabrication parameters found above, we fabricated a set of textile pneumatic actuators. We assess the performance of these actuators in force transmission and in cyclical testing. We choose these tests to evaluate the ability to produce normal force, to compare to similar pneumatic actuators [duPasquierHaptiknit:Haptics], and to ensure that the chosen fabrication method is robust enough to be used for wearable haptic devices, in which each actuator will undergo thousands of cycles. The force characterization tests (n=5) are conducted following a procedure similar to that outlined in [duPasquierHaptiknit:Haptics] and adapted for textile pneumatic actuators. As shown in Fig. 4A, the actuator is placed on a 3D-printed platform with a hole in it for the actuator to expand through. When inflated, the actuator presses against a force sensor (ATI Nano17 6-axis Force/Torque Sensor) below. Above the actuator, an acrylic plate prevents the actuator from expanding away from the force sensor. The platform, force sensor, and acrylic plate are secured to a 3D printed frame. The actuator is inflated by a pneumatic controller similar to the one proposed by du Pasquier et al. [duPasquierHaptiknit:Haptics]. During the test, the actuator is inflated to a series of absolute pressures from 125 to 230 kPa at 15 kPa intervals (a total of 8 peaks) and vacuumed to atmospheric pressure in between each pressure. The force is recorded at 100 Hz by the force sensor.
Cyclical testing (n=4) is performed with the same frame and load cell setup. Actuators are inflated to three target pressures, 150, 175, and 200 kPa, for 400 cycles. During each cycle, the actuator is inflated to the target pressure, held for one second, vacuumed to atmospheric pressure, and held for one second. Again, the force sensor records the output force at a frequency of 100 Hz. For each cycle, one peak force value is extracted from the data.
@@ III-B Results
Compared to the 3D-printed elastomeric pneumatic actuators from [duPasquierHaptiknit:Haptics], these textile pneumatic actuators have higher consistency in the cycling tests. At 150 and 200 kPa, the output force of the elastomeric pneumatic actuators drifted by a few Newtons over time. This is likely caused by the elastomeric actuators stretching and deforming with repeated use.
@@ IV Discussion and Conclusions
We investigated the effects of several fabrication parameters on the process of creating heat-sealed textile pneumatic actuators: temperature, pressure, and pressing duration. We conclude that the best parameters for fabricating the actuators are to heat press the TPU-coated nylon at 260 °F for 90 s under 5 MPa of pressure. These parameters are specific to our setup. Different results may be obtained by varying factors such as the type or thickness of the textile or by using a heat press with a single heated plate instead of two. Independent of exact values, we show that temperature, time, and pressure can have significant effects on the peel strength, peak force, and critical displacement of heat-sealed textiles, and outline a benchmarking method to evaluate the best fabrication parameters for any given setup. We also find that Gear Aid adhesive creates the best bond to our materials, and we can conclude more broadly that key features in the choice of an adhesive for textile pneumatic actuators are the following: repeatability, packaging, and post-cure adhesion, smoothness, and pliability.
@@
In testing, our actuators produce up to 36.1 N at 230 kPa absolute pressure when placed in a blocked force testing setup, which is 95.3% of the force produced by elastomeric pneumatic actuators of a comparable size in a similar setup [duPasquierHaptiknit:Haptics]. In contrast, textile pneumatic actuators achieve a 96.4% reduction in thickness when uninflated, decreasing the profile from 8.3 mm to 0.3 mm. The textile pneumatic actuators also achieve a 57.2% reduction in mass, from 4.73 g to 2.03 g. This allows our textile pneumatic actuators to be incorporated into lower-profile wearable devices without significantly compromising performance. Additionally, the textile actuators demonstrated more consistent performance over time during cycling tests, suggesting greater reliability for long-term use.
@@
Jumet et al. pioneered similar textile pneumatic actuators in bracelets [Jumet2022ADevices] and a sleeve [Jumet2023FluidicallyTextiles]. For these devices, a series of actuators were produced by heat-pressing two sheets of TPU-coated nylon taffeta fabric, between which a laser-cut paper layer prevents bonding in the inflation areas. This paper layer is not removable, but it allows for more convenient customization in actuator size and shape, as well as patch repairs in the case of punctured actuators. Like the actuators we tested, these actuators are compliant, comfortable, and low-profile. Their 25 mm square cells offer a 625 mm2 contact area, while our 25 mm round actuators offer a 491 mm2 contact area. At 100 kPa (200 kPa absolute pressure), their actuators produce up to 9.8 N of output force in a test setup mimicking the stiffness of the human wrist. In this setup, some of the force is absorbed deforming the synthetic skin, and the actuator may expand away from the force sensor, resulting in a lower measured force than a blocked-force test. In our blocked-force setup, we measured up to 30.6 N at the same pressure, but without losses. Assuming skin deformation and outward expansion reduce the force output by no more than 68%, our force measurements indicate that our actuator geometry is more effective at converting air pressure into force. These results also highlight the value of distributed-stiffness textiles, as described in [duPasquierHaptiknit:Haptics], for controlling the deformation of embedded pneumatic actuators. Layering stiff textiles over the actuator mimics a blocked-force setup by directing actuator deformation in a single direction.
Figure 5: Modular textile actuators integrated into a portable vest for haptic feedback on the torso.
@@
Due to to their low profile and rapid and repeatable fabrication method, textile pneumatic actuators can be integrated into a variety of wearable devices, such as bracelets, sleeves, collars, pants, or vests, as illustrated in Fig. 5. Their form factor allows for convenient and discrete integration into garments, increasing their wearability and user appeal. The soft textile surface of the actuators allows for comfortable nonadhesive direct-to-skin contact. Their compact inflation volume allows for operation with a portable controller equipped with an onboard pump and vacuum, thereby enhancing the potential for untethered wearability. Potential applications for pneumatic haptic garments with include improving immersion in virtual reality [Delazio2018ForceExperiences][Jadhav2023ScalableActuators][Kang2023PneumaticSimulation], replicating social touch in digital environments [Talhan2024SoftTouch][Keng2008HuggySystem], assisting navigation [Raitor2017WRAP:Guidance][Jumet2023FluidicallyTextiles], or providing guidance and feedback for posture or breathing [Wu2010Torso-basedDisorders][Rotella2012HAPIInterface].